OpenStudy (anonymous):

How do you use the square root method for the problem (y+7)^2=64

OpenStudy (anonymous):

y^2+14Y+49=64

OpenStudy (anonymous):

Take the square root of both sides \[y+7=\pm \sqrt{64}\]

OpenStudy (anonymous):

y+7=8

OpenStudy (anonymous):

Its plus or minus... so you have two equations y+7=8 and y+7=-8 y=1 and y=-15